OpenStudy (anonymous):

Solve the proportion 2 times x all over 8 times x minus 7 equals 2 over x.. What is the value of x? Answer x = -1 and x = -7 x = -1 and x = 8 x = 1 and x = -8 x = 1 and x = 7

sam (.sam.):

\[\frac{2x}{8x-7}=\frac{2}{x}\] \[x=7\]

sam (.sam.):

x = 1 and x = 7